# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

### v1

@@ 5 Experiments
To evaluate the effectiveness of TongUI agent, we conduct experiments on multiple GUI benchmarks in both offline and online settings.
@@ 5.1 Setting
Additional Training Data. In addition to GUI-Net, we also refine some publicly available GUI datasets. For GUI grounding, we adopt ShowUI-website [17], ShowUI-desktop [17], and AMEX [2]. For GUI navigation, following the guidance of ShowUI, we use the training set from AITW [29], Mind2Web [5], MiniWob [11], and GUIAct [3]. We prompt GPT-4o-mini to generate thoughts for these datasets with textual guidance and visual spots on images, obtaining 256K data. We combine these data with GUI-Net for training, totally 143K+256K=399K data.
Evaluation Data. We evaluate the TongUI agent on offline benchmarks: ScreenSpot [4], AITW [29], Mind2Web [5], and Baidu Experience (we manually annotate 102 data points from the collected Baidu Experience data for testing). We also evaluate the TongUI agent on one online benchmark: MiniWob [33]. We use the same 35 tasks selected by SeeClick[4] to evaluate our model.
Training. We use Qwen2.5-VL-3B/7B as our base vision-language-action model. We train our model with a context window of 8192. We set the maximum number of previous observations as 2, which results in at most three images in a trajectory. We limit each image to 1350 vision tokens. Our training involves two stages. First, we pre-train the model with datasets from ShowUI-desktop[17], ShowUI-web[17], AMEX[2], GUIAct[3], and GUI-Net. Secondly, we utilize the SFT data (training set of the evaluation dataset) to fine-tune the model. In both stages, we apply LoRA on our model’s vision encoder and LLM parts with rank 16 and alpha 32. This setting leads to 0.5% of total parameters. We set the batch size per device to 4 and the gradient accumulation steps to 8. We set the learning rate as $41e-$ 4 and AdamW optimizer for training. The pre-training process on 8$\times$×A100-80G takes roughly two days for a 3B model and three days for a 7B model.
Metrics. After training, we perform zero-shot evaluation of the model’s grounding capability on ScreenSpot. The annotation of ScreenSpot gives a bounding box for each interactive element. The action from the model is considered correct when the point predicted falls within the bounding box. For online evaluations, the environment of MiniWob provides reward functions, which suggests if a trajectory is correct. For offline evaluations such as Mind2Web, AITW, and Baidu Experience, the model will be asked to output actions with parameters. We evaluate if the action is correct by the exact match. For parameters, we do the exact match for string parameters and check if a point parameter falls within the bounding box based on annotations. We evaluate the model after fine-tuning on Mind2Web, AITW, and MiniWob. In other evaluation datasets, we use the pre-trained model.
Compared Methods. We compare the TongUI agent with open-source methods, including ShowUI [17], SeeClick [4], and UI-TARS [27]. The training data size (counted in tokens) for TongUI, ShowUI, and UI-TARS are 0.72B, 1.24B, and 50B, respectively. Our training set is much smaller than that of UI-TARS. Note that the data used in UI-TARS has not been released, and the number can be inferred from its technical report. We will release our data.
@@ 5.2.1 Grounding
In Tab. 1, we show the zero-shot grounding performance of TongUI. Similar to previous works [17], grounding on icon is much harder than grounding on text. The collected data leads to significant improvements on the baseline Qwen2.5-VL model. Compared with ShowUI, we use similar training data size, but it has about 4.5 % improvements. TongUI has a slightly worse performance than UI-TARS (3% on the 3B model and 6% on the 7B model), but the training data used by UI-TARS is 40 times that of ours, causing more resource consumption. These results demonstrate that GUI-Net can indeed improve the grounding capability of GUI agents.
@@ 5.2.2 Offline Navigation
TongUI-3B 48.0 88.4 44.2 48.9 85.4 42.6 50.0 87.7 46.0
TongUI-7B 51.1 88.7 46.9 50.4 87.5 43.7 53.9 88.6 49.1
Table 3: Results on Mind2Web. We report results on three types of tasks: cross-task, cross-website, and cross-domain. ‘Elem. Acc’ means whether the element is selected correctly, ‘OP. F1’ denotes the F1 score for the predicted action, and ‘Step SR’ counts successful steps.
Baidu Experience. The collected Baidu experience evaluation set includes tasks in Chinese, which evaluate the performance of GUI agents across different languages. Results are shown in Tab. 4. Compared with the baseline model, TongUI has improvements on both action accuracy and success rate, showing better generalization capability brought by the GUI-Net dataset.
Model Action Accuracy Success Rate
@@
Table 4: Results on the Baidu Experience
@@ 5.2.3 Online Navigation
TongUI-3B $\checkmark$ 71.7
TongUI-7B $\checkmark$ 71.9
Table 5: Results on the MiniWob
@@ 5.3 User Study for Data Quality
To evaluate the data quality of the collected GUI-Net dataset, we conduct a user study. Concretely, we randomly sample 100 data points from the GUI-Net dataset and 100 data points from the ShowUI dataset. We mix the 200 data points. We ask 5 persons (with rich experience in GUI agents but not familiar with our project) to provide scores (0-5) for the data quality, and they do not know whether the data is from GUI-Net or ShowUI. ‘0’ represents ‘very poor’ and ‘5’ represents ‘excellent’.
The results are shown in Fig. 5, where the average scores assigned to the data points from the ShowUI dataset and our GUI-Net dataset are calculated. The ShowUI dataset received an average score of 4.26, while the GUI-Net dataset received an average score of 4.12. This suggests that data quality in GUI-Net is comparable to that in the ShowUI dataset, while our data leads to further improvements.
@@
5.4 Data Scaling Effect
We evaluate the scaling law of GUI agents using collected data from different sources. Concretely, (1) we report the performance w/o SFT; (2) we use the refined public data to tune the model; (2) we add data collected from WikiHow to tune the model; (4) we add data collected from Baidu Experience to tune the model; (5) we add data collected from Video tutorials to tune the model. Results is shown in Tab. 6. We find that with the increase of data numbers and data sources, the performance of GUI agents is also constantly improving.
Data ScreenSpot Mind2Web
Task Website Domain
No SFT 8.0 0.4 1.0 1.7
Refined data 68.0 39.7 35.5 40.7
+ WikiHow 75.8 42.1 39.6 44.4
+ Baidu 78.7 43.4 41.6 45.5
+ Video 79.6 44.2 42.6 46.0
Table 6: Performance scaling over data sources testing on Qwen2.5-VL-3B, where Step SR is reported for Mind2Web.
@@ 6 Conclusion
In this paper, we present the TongUI framework that allows GUI agents to learn from multimodal web tutorials. By defining suitable and rich sources, we can crawl multimodal tutorials about diverse applications in different operating systems. The proposed tutorial processing method can extract tasks and trajectories from tutorials, and we obtain the GUI-Net dataset. Using this dataset, the tuned Qwen2.5-VL model achieves improvements on multiple commonly used benchmarks, demonstrating the effectiveness of the TongUI framework and the collected GUI-Net dataset. In this method, we have to collect all tutotrials and train the model once, which may overlook the potential for continual learning. In the future, we will explore the continual learning capability of GUI agents, enabling the agents to better adapt to new environments.
Appendix A Details
A.1 Action Space
The used action space is shown in Tab. 7.
Environment Action Definition
Desktop Click(x,y) Left single click on an element.
Input(value, x, y) Type a string into certain element.
Scroll(direction, x, y) Scroll screen with left/right/up/down direction. X and y is the start point.
LeftClickDouble(x, y) Left double click on an element
RightClickSingle(x, y) Right single click on an element
Drag(x1,y1,x2, y2) Drag the cursor to the position with left key pressed.
HotKey(value) Press the hot key.
Hover(x, y) Hover over an element.
Finish() Finish the task.
Mobile Tap(x, y) Finger tap on an element.
Input(value, x, y) Type a string into an element.
Swipe(x1, y1, x2, y2) Swipe the screen from the start position to the end position.
LongPress(duration, x, y) Press on an element with certain seconds.
PressHome() Press the home button.
PressBack() Press the back button.
Finish() Finish the task.
Table 7: Action space in trajectory generation.
A.2 Prompt
In LABEL:lst:systemPrompt, we show the system prompt in our agent framework.
Listing 1: System prompt for GUI-Net Agent
You are an assistant trained to navigate the computer/mobile screen.
Given a task instruction, a screen observation, and an action history sequence,
output the next action and wait for the next observation.
Here is the action space:
{action_space}
Format your response as
Thought: <your reasoning process>
Action: <the next action>
Format the action as a JSON object with the following keys:
{"action": "ACTION_TYPE", "value": "element", "position": [x,y]}
You can output multiple actions at once, and use JSON array to represent multiple actions.
If value or position is not applicable, set it as None‘.
Position might be [[x1,y1], [x2,y2]] if the action requires a start and end position.
Position represents the relative coordinates on the screenshot and should be scaled to a range of 0-1.
In LABEL:lst:actionSpace, we show a typical action space prompt for desktop.
Listing 2: Action space for GUI-Net Agent
1. CLICK‘: Click on an element, value is not applicable and the position [x,y] is required.
2. INPUT‘: Type a string into an element, value is a string to type and the position [x,y] is required.
3. SCROLL‘: Scroll the screen, value is the direction to scroll and the position is start position of the scroll operation.
4. LEFT_CLICK_DOUBLE‘: Left click on an element twice, value is not applicable and the position [x,y] is required.
5. RIGHT_CLICK_SINGLE‘: Right click on an element once, value is not applicable and the position [x,y] is required.
6. DRAG‘: Drag the cursor to the specified position with the left button pressed. Value is not applicable and position [[x1,y1], [x2,y2]] is the start and end position of the drag operation.
7. HOT_KEY‘: Press a hot key, value is the hot key and the position is not applicable.
8. WAIT‘: Wait for 5 seconds, and take a screenshot to check for any changes. Value and position are not applicable.
9. FINISH‘: Finish the task. Value and position are not applicable.
Appendix B Dataset
B.1 Dataset Size
We show our data size in Tab. 8, including four components: collected data from Wikihow, collected data from Baidu Experience, collected data from Video tutorials, and refined data from ShowUI.
Dataset Token Trajectory
Wikihow 151M 42K
Baidu Experience 178M 51K
Video Tutorial 195M 50K
Refined ShowUI 720M 256K
Table 8: Our data statistics by the token and trajectory. We compute token by Qwen2.5-VL tokenizer and each image is counted as 1350 vision tokens.
B.2 Applications
In GUI-Net, the collected data spans about 280 applications that are shown in Figs. 6 and 7.
Figure 6: Collected applications.
Figure 7: Collected applications.
Appendix C Examples
C.1 GUI-Net
We visualize several examples of collected trajectory data, as shown in Figs. 8, 9, 10, 11 and 12.
Figure 8: Dataset example.
Figure 9: Dataset example.
Figure 10: Dataset example.
Figure 11: Dataset example.
Figure 12: Dataset example.
C.2 TongUI
We visualize several examples executed by the TongUI agent, as shown in Figs. 13, 14, 15 and 16.
Figure 13: Examples executed by TongUI.
Figure 14: Examples executed by TongUI.
Figure 15: Examples executed by TongUI.
Figure 16: Examples executed by TongUI.

### v2

@@ 5.1 Setting
Additional Training Data. In addition to GUI-Net-1M, we also refine some publicly available GUI datasets. For GUI grounding, we adopt ShowUI-website [18], ShowUI-desktop [18], and AMEX [2]. For GUI navigation, following the guidance of ShowUI, we use the training set from AITW [31], Mind2Web [5], MiniWob [11], and GUIAct [3]. We prompt GPT-4o-mini to generate thoughts for these datasets with textual guidance and visual spots on images, obtaining 256K data. We combine these data with GUI-Net-1M for training, totally 1M+256K $\approx$ 1.3M data. To make sure the thought is consistent with the action, we explicitly add vision prompts to images, such as a spot that indicates the location of the clicked button. In our text prompt, we instruct the model to generate a thought based on textual guidance of the task, description of the action, and vision prompts on the image.
Evaluation Data. We evaluate the TongUI agent on offline benchmarks: ScreenSpot [4], ScreenSpot-V2 [39], ScreenSpot-Pro [15], AITW [31], AndroidControl [16], Mind2Web [5], and Baidu Experience (we manually annotate 102 data points from the collected Baidu Experience data for testing). We also evaluate the TongUI agent on an online benchmark: MiniWob [34]. We use the same 35 tasks selected by SeeClick[4] to evaluate our model.
Training. We use Qwen2.5-VL-3B/7B as our base vision-language-action model. We train our model with a context window of 8192. We set the maximum number of previous observations as 2, which results in at most three images in a trajectory. We limit each image to 1350 vision tokens. Our training involves two stages. Following existing methods [18], we utilize the SFT data (training set) of the evaluation dataset to fine-tune the model. We apply LoRA on our model’s vision encoder and LLM parts with rank 16 and alpha 32. This setting leads to 0.5% of total parameters. We set the batch size per device to 4 and the gradient accumulation steps to 8. We set the learning rate as $41e-$ 4 and AdamW optimizer for training. The training process on 8$\times$×A100-80G takes roughly four days for a 3B model and seven days for a 7B model.
Compared Methods. We mainly compare the TongUI agent with open-source methods, including ShowUI [18], SeeClick [4], OS-Atlas [39], CogAgent [10], etc.
Additional experimental results, such as data scaling effect, are presented in the Supplementary Material.
@@ 5.2.1 Grounding
In Tab. 1,  Tab. 2, and  Tab. 3, we show the zero-shot grounding performance of TongUI on ScreenSpot, ScreenSpot-V2, and ScreenSpot-pro, respectively. Similar to previous works [18], grounding on icon is much harder than grounding on text. The collected data leads to significant improvements on the baseline Qwen2.5-VL model. Compared with ShowUI, our method has about 5 % - 20 % improvements. These results demonstrate that GUI-Net-1M can indeed improve the grounding capability of GUI agents.
@@ 5.2.2 Offline Navigation
Method Model High Low
AITW PaLM 2L 19.5 56.7
SeeClick GPT-4-Turbo 33.9 54.3
M3A GPT-4-Turbo 42.1 55.0
ER PALM-2S 19.5 45.5
ER PALM 2L 33.0 45.9
ER GPT-4 32.1 51.7
ER Gemini 1.5 Pro 24.4 50.2
TongUI TongUI-3B 32.0 52.0
TongUI TongUI-7B 38.2 69.3
Table 5: Step accuracy on AndroidControl.
AndroidControl. We also evaluate our performance in the AndroidControl benchmark in Tab. 5. As we do not include the AndroidControl training set, we compare our model with the zero-shot methods mentioned in previous work  [16]. TongUI-7B model achieves the best performance of low-level instructions(12.6% improvement).
@@
TongUI-3B 53.4 89.0 48.8 54.2 86.4 48.1 53.8 88.2 49.5
TongUI-7B 58.1 88.7 53.4 55.6 87.2 49.0 57.6 88.7 52.9
Table 6: Results on Mind2Web. We report results on three types of tasks: cross-task, cross-website, and cross-domain. “Elem. Acc” means whether the element is selected correctly, “OP. F1” denotes the F1 score for the predicted action, and “Step SR” counts successful steps.
Baidu Experience. The collected Baidu experience evaluation set includes tasks in Chinese, which evaluate the performance of GUI agents across different languages. Results are shown in Tab. 7. Compared with the baseline model, TongUI has improvements on both action accuracy and success rate, showing better generalization capability brought by the GUI-Net-1M dataset.
Model Action Acc. Success
@@
Table 7: Results on Baidu Experience.
@@ 5.2.3 Online Navigation
TongUI-3B $\checkmark$ 72.7
TongUI-7B $\checkmark$ 73.9
Table 8: Results on MiniWob.
@@ 5.3 User Study for Data Quality
To validate the effectiveness of our data filtering strategy and evaluate the data quality of the collected GUI-Net-1M dataset, we conduct two user studies. In both studies, five participants with substantial experience in GUI agent research (but not involved in our project) are asked to rate the quality of each data point on a scale from 0 (very poor) to 5 (excellent). As shown in Fig. 6(a), we evaluate the impact of our Trajectory Filtering step. Specifically, we randomly sample 100 data points before filtering and 100 after filtering, and mix them. As shown in Fig. 6(a), the average score increases from 3.22 to 4.12 after filtering, demonstrating that Trajectory Filtering significantly improves data quality by effectively removing low-quality or GUI-irrelevant steps. We also randomly sample 100 data points from the ShowUI dataset for comparison. The results are shown in Fig. 6(b), the ShowUI dataset received an average score of 4.26. This suggests that data quality in GUI-Net-1M is comparable to that in the ShowUI dataset, while our data leads to further improvements.
@@ 6 Conclusion
In this paper, we present the TongUI framework that allows GUI agents to learn from multimodal web tutorials. By defining suitable and rich sources, we can crawl multimodal tutorials about diverse applications in different operating systems. The proposed tutorial processing method can extract tasks and trajectories from tutorials, and we obtain the GUI-Net-1M dataset. Using this dataset, the tuned Qwen2.5-VL model achieves improvements on multiple commonly used benchmarks, demonstrating the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset. In this method, we have to collect all tutotrials and train the model once, which may overlook the potential for continual learning. In the future, we will explore the continual learning capability of GUI agents, enabling the agents to better adapt to new environments.